7/(2x-12) + 2 = 1/(x-6)

Answer by ajeet(1)

7/2(x-6) +2 = 1/(x-6)
1/(x-6)-7/2(x-6)= 2
-5/2(x-6)=2
-5 = 4(x-6)
-5 = 4x-24
4x = -5+24 = 19
x = 19/4